Use the quadratic formula to solve each equation please :)
(x+1)(x-2)=5



x^2+9x-2=-16

Answer by addingup(3677)

(x+1)(x-2) = 5; x^2-2x+1x-2 = 5; x^2-1x-2
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
x^2+9x-2=-16: cannot factor